Find the​ mean, variance, and standard deviation of the binomial distribution with the given values of n and p. n equals 70 ​, p equals 0.3 The​ mean, mu ​, is nothing . ​(Round to the nearest tenth as​ needed.)
Answer #1


Given
n = 70
p = 0.3

We know that for a binomial distribution

Mean
mu = n imes p = 70 imes 0.3 = 21
Variance
σ2 = n × p × (1-p) = 70 × 0.3 × 0.7 14.7
Standard deviation
sigma = sqrt{n imes p imes(1-p)} = sqrt{14.7}=3.83